# Statistical Validation and Vetting of Exoplanet Candidate TOI 7475.01

### v1

@@ Abstract
We present a comprehensive validation analysis of the exoplanet candidate TOI 7475.01 (TIC 376866659), detected by the TESS mission. Using a custom pipeline combining natural flux preservation with robust BLS detection, we identified a transit signal with a period of $3.2538$ days and a depth of $\sim 4600$ ppm. To rule out false positives, we performed centroid analysis, spatial contamination checks using Gaia DR3, and a statistical validation using triceratops. Our results show a Signal-to-Noise Ratio (SNR) of 294.13 and a False Positive Probability (FPP) of $\approx 0$. Based on the clean spatial environment, stable centroids, and high statistical probability, we validate TOI 7475.01 as a planetary companion.
exoplanets — TESS — techniques: photometric — methods: statistical — stars: individual (TIC 376866659)
@@ I Introduction
The Transiting Exoplanet Survey Satellite (TESS) has revolutionized the search for exoplanets around nearby bright stars (Ricker et al., 2015). In this work, we focus on TOI 7475.01, a candidate listed on the Exoplanet Follow-up Observing Program (ExoFOP) website (Akeson et al., 2013), associated with TIC 376866659. The goal of this study is to vet the signal against common astrophysical false positives, such as eclipsing binaries (EBs) and background contamination.
@@ IV.2 Scenario Probability Breakdown
  • TP (Transiting Planet): $\sim 45\%$. The planet orbits the target star. This is the most likely single scenario.
  • PTP (Primary Transiting Planet): $\sim 40\%$. The planet orbits an unresolved bound companion. In this scenario, the planet still belongs to the target system.
  • DTP (Double Transiting Planet): $\sim 15\%$. An unresolved background star is present in the aperture, but the transiting planet orbits the target star. Although the presence of an unresolved background star seems unlikely given the clean Gaia DR3 data (Section 3.1), even in this scenario, the transit occurs on the target.
@@
V Conclusion and Future Work
@@
Based on the high SNR detection (294.13), the clean spatial environment, stable centroid analysis, and a False Positive Probability of FPP $\approx 0$, we conclude that TOI 7475.01 is a statistically validated exoplanet.
The signal exhibits all the characteristics of a genuine planetary transit, with no evidence of contamination from background sources or binary stars.
Future Work (v2)
While this work confirms the planetary nature of the candidate, precise determination of the planetary radius and orbital inclination requires more sophisticated modeling.
In the upcoming Version 2 (v2) of this paper, we intend to perform a full planetary characterization using the juliet library (Espinoza et al., 2019). This will involve applying Nested Sampling algorithms to derive robust posterior distributions of the planetary parameters and modeling the stellar density to better constrain the impact parameter and eccentricity.
@@
The TESS data presented in this paper can be obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI). The specific data products used (Sector 91) are publicly available under the target TIC 376866659.
@@ Acknowledgments
This research made use of Lightkurve, a Python package for Kepler and TESS data analysis; TRICERATOPS for statistical validation; Astropy, a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2022); NumPy (Harris et al., 2020); Matplotlib (Hunter, 2007); and pandas.

### v2

# Statistical Validation and Photometric Characterization of the Hot Jupiter Candidate TOI 7475.01

@@ Abstract
We present the statistical validation and full photometric characterization of the exoplanet candidate TOI 7475.01 (TIC 376866659), detected by the TESS mission. Using a custom pipeline combining natural flux preservation with robust BLS detection, we identified a transit signal with a period of $3.2538$ days and a depth of $\sim 4600$ ppm. To rule out false positives, we performed centroid analysis, spatial contamination checks using Gaia DR3, and a statistical validation using TRICERATOPS. Our results show a Signal-to-Noise Ratio (SNR) of 294.13 and a False Positive Probability (FPP) of $\approx 0$. Based on the clean spatial environment, stable centroids, and high statistical probability, we validate TOI 7475.01 as a planetary companion. We subsequently performed a Bayesian transit fit using juliet with Dynesty nested sampling, and derived physical parameters via Monte Carlo error propagation. We find a planetary radius of $R_{p}=1.18^{+0.39}_{-0.40}\ R_{\mathrm{Jup}}$ and an equilibrium temperature of $T_{\mathrm{eq}}=1455^{+77}_{-56}$ K, consistent with a hot Jupiter classification. The planet mass is estimated at $M_{p}\approx 2.2\ M_{\mathrm{Jup}}$ (MAP) via the Chen & Kipping (2017) mass-radius relation; radial velocity follow-up is required for a precise measurement. The impact parameter remains unconstrained ($b=0.46\pm 0.34$), a limitation of single-band TESS photometry that future CHEOPS observations could resolve.
exoplanets — TESS — hot Jupiters — techniques: photometric — methods: statistical — methods: Bayesian — stars: individual (TIC 376866659)
@@ I Introduction
The Transiting Exoplanet Survey Satellite (TESS) has revolutionized the search for exoplanets around nearby bright stars (Ricker et al., 2015). In this work, we focus on TOI 7475.01, a candidate listed on the Exoplanet Follow-up Observing Program (ExoFOP) website (Akeson et al., 2013), associated with TIC 376866659.
This paper combines two stages of analysis: the statistical validation of the candidate (Sections IIIV), which confirms the planetary nature of the signal, and a full photometric characterization (Sections VVII), which derives the physical properties of the planet via Bayesian modeling. The goal of the first stage is to vet the signal against common astrophysical false positives, such as eclipsing binaries (EBs) and background contamination. The goal of the second stage is to determine the orbital and physical parameters of the validated companion.
@@ IV.2 Scenario Probability Breakdown
  • TP: $\sim 45\%$. The planet orbits the target star. This is the most likely single scenario.
  • PTP: $\sim 40\%$. The planet orbits an unresolved bound companion. In this scenario, the planet still belongs to the target system.
  • DTP: $\sim 15\%$. An unresolved background star is present in the aperture, but the transiting planet orbits the target star. Although the presence of an unresolved background star seems unlikely given the clean Gaia DR3 data (Section III), even in this scenario, the transit occurs on the target.
@@
V Bayesian Transit Fit
V.1 Light Curve Preprocessing
For the Bayesian fit we applied transit windowing, retaining only data within $1.5\times T_{14}$ of each predicted transit center, reducing the dataset to 1,641 binned data points from the original 86,906-point SPOC light curve. The pre-fit flux diagnostic confirms robust transit detection: the observed in-transit flux decrement is 4735-4735 ppm, consistent with the BLS depth of 4601-4601 ppm (3% discrepancy, within noise). Figure 5 shows the phase-folded light curve with 10-minute binning.
Figure 5: Pre-fit Phase-folded Light Curve. Top: zoom around the transit. Black points are 10-minute bins; the red shaded region marks the expected transit window; the dashed line marks the expected depth. Bottom: full phase view showing the out-of-transit baseline. The transit is clearly detected at the expected depth and epoch.
V.2 Model and Priors
We performed a Bayesian transit fit using juliet (Espinoza et al., 2019), which interfaces with the Dynesty nested sampler (Speagle, 2020). We adopted a quadratic limb darkening law in the Kipping (2013) parametrization, with Gaussian priors on the coefficients centered on the Claret (2017) theoretical values for the stellar parameters of TIC 376866659: $q_{1}=0.348\pm 0.05$, $q_{2}=0.246\pm 0.05$.
The eccentricity was fixed to zero, justified by the short orbital period ($P=3.25$ days) at which tidal circularization timescales are expected to be shorter than the system age (Goldreich & Soter, 1966).
The stellar mean density $\rho_{*}=350.9\pm 93.1$ kg m-3 (computed from $M_{*}$ and $R_{*}$) was included as a free parameter with a Gaussian prior. This allows juliet to compute $a/R_{*}$ self-consistently via Kepler’s third law reformulated in terms of density:
$\frac{a}{R_{*}}=\left(\frac{G\,\rho_{*}\,P^{2}}{3\pi}\right)^{1/3},$ (1)
which links the transit light curve directly to a stellar observable independent of the planet.
The prior on the radius ratio $p=R_{p}/R_{*}$ was set uniform over $[0.034,0.102]$, bracketing $0.5\times$ to $1.5\times$ the value expected from the BLS depth ($p_{\mathrm{exp}}=\sqrt{\delta}=0.068$). All prior settings are summarized in Table 3.
The nested sampling was run with 1,000 live points and convergence criterion $\Delta\log Z<0.1$, requiring approximately 323,000 likelihood evaluations.
Table 3: Prior Distributions for the Bayesian Fit
Parameter Prior Value
$P$ [days] Normal $3.253773\pm 0.0001$
$T_{0}$ [BTJD] Normal $3775.582\pm 0.005$
$p=R_{p}/R_{*}$ Uniform [0.034, 0.102][0.034,\ 0.102]
$b$ Uniform [0.0, 0.99][0.0,\ 0.99]
$\rho_{*}$ [kg/m³] Normal $350.9\pm 93.1$
$e$ Fixed 0
$q_{1}$ Normal $0.348\pm 0.05$
$q_{2}$ Normal $0.246\pm 0.05$
$\sigma_{w}$ [ppm] Log-unif. [0.1, 1000][0.1,\ 1000]
V.3 Posterior Constraints
Figure 6 shows the corner plot of the fit posteriors. The period and epoch are well-constrained ($\sigma_{P}\sim 10^{-4}$ days, $\sigma_{T_{0}}\sim 5\times 10^{-3}$ days), recovering the BLS solution with minimal improvement, as expected given the tight Gaussian priors.
The radius ratio converges to $p=0.07^{+0.02}_{-0.02}$. In contrast, the impact parameter $b=0.46^{+0.36}_{-0.32}$ remains essentially flat across its entire prior range $[0,0.99]$, indicating that the TESS data cannot distinguish between a central and a grazing transit geometry. This ppbb degeneracy arises because a larger planet on a grazing orbit produces the same integrated flux decrement as a smaller planet on a central orbit; resolving it requires accurate measurement of the ingress and egress durations, which demands higher cadence or higher photometric precision than available here. As a direct consequence, the stellar density posterior ($358\pm 93$ kg m-3) reproduces the prior almost exactly, since $\rho_{*}$ is constrained by $a/R_{*}$ which itself depends on $b$.
Figure 6: Corner Plot of Fit Posteriors. Diagonal panels: marginalized 1D histograms with the 16th, 50th, and 84th percentiles marked. Off-diagonal panels: 2D joint distributions. Note the broad, essentially flat posterior of $b$ (second column/row) and its correlation with $p=R_{p}/R_{*}$, which limits the precision of the radius determination.
VI Planetary Parameters
VI.1 Monte Carlo Error Propagation
Physical parameters were derived via Monte Carlo error propagation with N=10, [^0] N=10{,}000 samples. At each iteration we drew independently from Gaussian distributions for $M_{*}$, $R_{*}$, and $T_{\mathrm{eff}}$, and from the Juliet posterior samples for $p$, $b$, $\rho_{*}$, $P$, and $T_{0}$. This approach correctly propagates the non-Gaussian shape of the posteriors into the derived quantities, avoiding the small-error assumptions of standard analytical propagation. In particular, it accounts for the simultaneous uncertainty in the fit (photometric data) and in the stellar parameters (spectroscopic observations), which would be impossible to combine analytically given their non-linear coupling.
The planet radius follows from $R_{p}=p\cdot R_{*}$, converted to Jupiter units using the equatorial ratio $R_{\odot}/R_{\mathrm{Jup}}=9.9549$. The semi-major axis is computed from Equation (1). The equilibrium temperature assumes Bond albedo $A=0.3$ and efficient day–night heat redistribution ($f=0.5$):
$T_{\mathrm{eq}}=T_{\mathrm{eff}}\left(\frac{R_{*}}{2a}\right)^{1/2}\!\left[f(1-A)\right]^{1/4}.$ (2)
For asymmetric posterior distributions, particularly $M_{p}$, we report both the median with 16th/84th percentile errors and the Maximum A Posteriori (MAP) value, estimated via Gaussian Kernel Density Estimation. The MAP is more representative of the most probable value when the distribution is strongly skewed.
VI.2 Planet Mass
In the absence of radial velocity measurements, the planet mass was estimated using the probabilistic mass-radius relation of Chen & Kipping (2017), implemented via the mr_forecast package. The resulting posterior is highly asymmetric: the bulk of the probability is concentrated at $M_{p}\sim 1$$3\ M_{\mathrm{Jup}}$, but the distribution extends to $\sim 200\ M_{\mathrm{Jup}}$ due to the degeneracy in the giant planet regime, where similar radii can correspond to a wide range of masses from standard gas giants to brown dwarfs. The MAP estimate ($M_{p}\approx 2.2\ M_{\mathrm{Jup}}$) is therefore more informative than the median ($3.1^{+64.7}_{-2.9}\ M_{\mathrm{Jup}}$), and should be interpreted as an order-of-magnitude estimate pending RV confirmation.
VI.3 Summary of Parameters
The complete set of derived parameters is given in Table 4. The posterior distributions for the main physical quantities are shown in Figure 7.
Table 4: Stellar and Planetary Parameters of TOI 7475.01
Parameter Symbol Value Source
Stellar parameters
Effective temperature $T_{\mathrm{eff}}$ [K] $6467\pm 52$ Spectroscopy
Stellar mass $M_{*}\ [M_{\odot}]$ $1.279\pm 0.200$ Spectroscopy
Stellar radius $R_{*}\ [R_{\odot}]$ $1.726\pm 0.100$ Spectroscopy
Surface gravity $\log g$ [cgs] $4.14\pm 0.10$ Spectroscopy
Metallicity $\mathrm{Fe/H}]$ $0.10\pm 0.08$ Spectroscopy
$J$-band magnitude $m_{J}$ $7.676$ 2MASS
Orbital parameters (Juliet posteriors)
Orbital period $P$ [days] $3.25377^{+0.00010}_{-0.00009}$ This work
Transit epoch $T_{0}$ [BJD] $2460775.5816^{+0.0050}_{-0.0050}$ This work
Eccentricity $e$ 0 (fixed) This work
Radius ratio $p=R_{p}/R_{*}$ $0.07^{+0.02}_{-0.02}$ This work
Impact parameter $b$ $0.46^{+0.36}_{-0.32}$ This work
Planetary parameters (Monte Carlo)
Planet radius (median) $R_{p}\ [R_{\mathrm{Jup}}]$ $1.18^{+0.39}_{-0.40}$ This work
Planet radius (MAP) $R_{p}\ [R_{\mathrm{Jup}}]$ $1.18$ This work
Planet mass (MAP) $M_{p}\ [M_{\mathrm{Jup}}]$ $\approx 2.2$ Chen & Kipping (2017)
Semi-major axis $a$ [AU] $0.047^{+0.005}_{-0.005}$ This work
Equilibrium temperature $T_{\mathrm{eq}}$ [K] $1455^{+77}_{-56}$ This work
Incident flux $F_{p}\ [F_{\oplus}]$ $2137^{+488}_{-309}$ This work
Planet density (MAP) $\rho_{p}$ [kg m-3] $\approx 1482$ This work
Atm. scale height (MAP) $H$ [km] $\approx 29$ This work
TSM ($M_{p}=1\ M_{\mathrm{Jup}}$) 17.1 Kempton et al. (2018)
Figure 7: Derived Posterior Distributions. Histograms of the six main physical parameters from the Monte Carlo error propagation (N=10, [^0] N=10{,}000 samples). Solid vertical lines: median. Dashed lines: 16th and 84th percentiles. The $M_{p}$ distribution (top center) is strongly asymmetric due to the mass-radius degeneracy in the giant planet regime; the MAP ($\approx 2.2\ M_{\mathrm{Jup}}$) is more representative than the median.
VII Discussion
VII.1 Classification
[…]
The MAP density ($\rho_{p}\approx 1482$ kg m-3) is consistent with a non-inflated hot Jupiter ($\rho_{\mathrm{Jup}}=1326$ kg m-3). This is noteworthy given the high irradiation ($F_{p}\approx 2137\ F_{\oplus}$), as ohmic heating and other inflation mechanisms are expected to be active at this flux level (Demory & Seager, 2011). However, the mass uncertainty is sufficiently large that this conclusion is not firm: a lower actual mass would imply a lower density and a more inflated structure.
VII.2 Population Context
Figure 8 places TOI 7475.01 in the context of 676 hot Jupiters ($P<10$ days, $R_{p}>0.5\ R_{\mathrm{Jup}}$, $T_{\mathrm{eq}}>500$ K) from the NASA Exoplanet Archive. The planet occupies a well-populated region at intermediate temperature ($\sim 1450$ K) and typical Jupiter radius, broadly consistent with the bulk of the known hot Jupiter population.
Figure 8: Hot Jupiter Population Comparison. Gray points: 676 hot Jupiters from the NASA Exoplanet Archive with known radii and equilibrium temperatures. Red star: TOI 7475.01 (median values). Dashed lines mark the inflated ($R_{p}>1.3\ R_{\mathrm{Jup}}$) and ultra-hot ($T_{\mathrm{eq}}>1800$ K) boundaries. TOI 7475.01 sits in a well-populated region at intermediate temperature and typical Jupiter size.
VII.3 Impact Parameter Degeneracy and Recommended Follow-up
The principal limitation of this characterization is the unconstrained impact parameter ($b=0.46^{+0.36}_{-0.32}$), which propagates into a large fractional radius uncertainty ($\sim 34\%$). This degeneracy is a known limitation of single-band TESS photometry and arises because the integrated transit depth is preserved across a family of $(p,b)$ solutions with the same $p^{2}$. As visible in the corner plot (Figure 6), the posterior of $b$ is essentially flat over its entire prior range, and the joint ppbb distribution shows no constraining power from the data alone.
High-precision follow-up photometry with CHEOPS (Benz et al., 2021) would break this degeneracy by resolving the ingress and egress shapes with sufficient precision to separately constrain $b$ and $p$. Ground-based multi-band photometry could additionally constrain the limb darkening independently, providing a further check on the stellar parameters.
[…]
VII.4 Atmospheric Observability
The Transmission Spectroscopy Metric (TSM, Kempton et al., 2018) for this system is TSM $\approx 17$ (assuming $M_{p}=1\ M_{\mathrm{Jup}}$), below the recommended threshold of 90 for hot Jupiter atmospheric characterization with JWST. The low TSM reflects the brightness of the host star in the near-infrared ($m_{J}=7.676$) combined with the moderate planet-to-star radius ratio. While the atmospheric scale height $H\approx 29$ km (MAP) is typical for hot Jupiters, this planet is a marginal JWST target compared to the best-characterized systems currently known.
VIII Conclusion
@@
Based on the high SNR detection (294.13), the clean spatial environment, stable centroid analysis, and a False Positive Probability of FPP $\approx 0$, we confirm that TOI 7475.01 is a statistically validated exoplanet. The signal exhibits all the characteristics of a genuine planetary transit, with no evidence of contamination from background sources or binary stars.
The subsequent Bayesian characterization yields the following main results:
  1. 1.
[…]
  2. 2.
    Equilibrium temperature $T_{\mathrm{eq}}=1455^{+77}_{-56}$ K, firmly classifying the system as a hot Jupiter.
  3. 3.
    Estimated mass $M_{p}\approx 2.2\ M_{\mathrm{Jup}}$ (MAP, Chen-Kipping), pending RV confirmation.
  4. 4.
    The impact parameter $b$ is unconstrained by the TESS data alone, limiting the precision of the radius determination. This is the principal observational limitation of this work.
  5. 5.
    CHEOPS follow-up photometry is recommended to break the ppbb degeneracy, and radial velocity observations are required to measure the true planetary mass.
@@
The TESS data presented in this paper can be obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute (STScI). The specific data products used (Sector 91) are publicly available under the target TIC 376866659. The analysis pipeline is publicly available at https://github.com/biesro/TESS-TOI-7475.01-Validation.
@@ Acknowledgments
This research made use of Lightkurve (Lightkurve Collaboration et al., 2018); TRICERATOPS (Giacalone et al., 2021); juliet (Espinoza et al., 2019); Dynesty (Speagle, 2020); mr_forecast (Chen & Kipping, 2017); corner.py (Foreman-Mackey, 2016); Astropy (Astropy Collaboration et al., 2022); NumPy (Harris et al., 2020); Matplotlib (Hunter, 2007); and pandas.